# To Be a Truster or Not to Be: Evolutionary Dynamics of a Symmetric N-Player Trust Game in Well-Mixed and Networked Populations

### v1

@@ Abstract
Evolutionary game theory, evolutionary dynamics, replicator dynamics, trust game, multiplayer game. symmetry, networks
@@ I-A Evolution of Prosocial Behaviours
Researchers across disciplines have explored how prosocial behaviours evolve among self-interested individuals. A key focus is the evolution of cooperation in social dilemmas, such as the Prisoner’s Dilemma (PD) game and its $N$-player variant, the Public Goods Game (PGG) [1, 2, 3, 4, 5, 6, 7, 8]. In particular, evolutionary game theory enables us to examine how successful strategies proliferate through evolution, i.e., fitness-dependent reproduction and imitation [9, 10]. Evolutionary game theory also has practical applications, including green supply chain policy formulation, modelling technology diffusion, and others [11, 12, 13].
Many real-world scenarios involve sequential interactions between two players, as seen in buyer-seller exchanges. Unlike the PD and PGG, which model simultaneous interactions, sequential interactions introduce a trust issue where one player’s decision may leave them vulnerable to exploitation by the other [14]. The concepts of trust and trustworthiness have also gained attention in engineering research [15, 16, 17, 18, 19]. Many problems in these fields are framed as buyer-seller interactions [20]. The Trust Game (TG) is a current standard for formalising non-simultaneous interactions under social dilemmas and is widely used for studying trust and trustworthiness [14, 21, 22, 23, 24, 25, 26, 27, 28]. The TG involves a one-shot sequential interaction between an investor (truster) and a trustee. The binary TG is a simple variant where each role has two strategies: investors can choose to invest or not, and trustees can choose to be trustworthy or untrustworthy [24, 29, 30, 31]. Evolutionary game theory predicts that anti-social strategies evolve in the two-player TG, resulting in reduced payoffs for both players compared to prosocial strategies. Consequently, an additional mechanism is needed for prosocial strategies to evolve in the TG [32, 33, 34]. In practice, interactions often involve multiple participants, so recent attention has been paid to $N$-player TG (NTG) [35, 36, 37, 38, 39, 40], which attempts to generalise the two-player TG to multiplayer settings, with $N\geq 3$.
I-B Previous Work on NTG
@@ I-B2 Asymmetric NTG
In both the original and asymmetric NTGs, single-role payoffs drive evolutionary game dynamics. As such, they may be less suitable for scenarios in which players alternate between the two roles (i.e., investor and trustee) and the average payoff over the two roles should drive the evolutionary dynamics. It is also worth noting that behavioural experiments often assign human participants to alternate between investor and trustee roles [47, 48, 49].
@@ I-C Contributions
We propose a symmetric NTG (SNTG) to model situations in which players alternate between the two roles. Our contributions include:
@@ II-A Strategies
Our SNTG expands upon the asymmetric NTG [39] by symmetrising it. Players alternate between investor and trustee roles, similar to symmetrising two-player asymmetric games [34, 52, 53]. Investors choose whether to invest or not, while trustees decide to be trustworthy or untrustworthy. A player’s strategy therefore comprises two elements, one for each role: $i$ (i.e., invest) and $n$ (i.e., not to invest) as investor and $t$ (i.e., trustworthy) and $u$ (i.e., untrustworthy) as trustee. There are four possible strategies: $it,iu,nt$, and $nu$.
@@ II-D2 Simulations
Unless we state otherwise, simulations commence with equal proportions of the four strategies ($it$, $iu$, $nt$, and $nu$) uniformly randomly distributed over the $Z=1024$ nodes, which we refer to as the unbiased initial condition. Each simulation runs for 5000 generations. In each generation, the two steps are repeated $Z$ times, allowing each node to update its strategy once per generation on average. For each parameter pair $(w,r)$, we ran 20 independent simulations.
@@ III-A Well-mixed Populations
Figure 3: Time courses of the SNTG in the infinite well-mixed population with initial conditions in the interior of state space $\triangle^{3}$. The fractions of the different strategies, $y_{it}(t)$, $y_{iu}(t)$, $y_{nt}(t)$, and $y_{nu}(t)$ are shown as a function of time for various initial conditions. (Top row) Unbiased initial state $y_{it}(0)=y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=1/4$. (Bottom) 256 randomly selected initial conditions. We only show $y_{it}(t)$ and $y_{iu}(t)$ for clarity. We observe $y_{it}(t)+y_{iu}(t)\rightarrow 0$ as $t\rightarrow\infty$, regardless of the initial condition; in other words, $y_{nt}(t)+y_{nu}(t)\rightarrow 1$ as $t\rightarrow\infty$. Hence, investment does not evolve. Parameters are the same as those used in Fig. 2.
In a well-mixed population, investment (i.e. trust) does not evolve (Figs. 2 and 3). Consequently, the average payoff for a player is equal to 00. All trajectories converge to the line of stable equilibria on the NT-NU edge, including the vertex NU (i.e., unanimity of $nu$ players), i.e., $y_{nt}+y_{nu}=1$ with $\frac{r}{r+1}<y_{nu}\leq 1$, of the 3-simplex $\triangle^{3}=\{(y_{it},y_{iu},y_{nt},y_{nu}):y_{it}+y_{iu}+y_{nt}+y_{nu}=1\}$ as the state space (Fig. 2).
Vertices IT, IU and NT of the 3-simplex are unstable equilibria. The entire interior of the IT-NU and IU-NT edges and part of interior of the NT-NU edge given by $0<y_{nu}<\frac{r}{r+1}$ are also unstable equilibria. There is no other equilibria including the interior of the triangles and the tetrahedron. For proofs, see Appendix A-F, A-G, and A-H. The nonlinearity in the payoff, $w$, has no impact on the evolutionary dynamics. Moreover, the evolutionary outcomes of the SNTG resemble those of the symmetric 2-player TG [34], with both games ultimately resulting in $y_{nt}+y_{nu}=1$, indicating no evolution of investment.
@@ III-B2 Heterogeneous Network
Figure 4: Properties of approximate equilibria of the SNTG in finite networks. (a) Square lattice. (b) Heterogeneous networks. In both (a) and (b), the first four columns show the fraction of each strategy as the function of $w$ and $r$. The fifth column shows the average payoff over all nodes, normalised by that of a population comprised entirely of $it$. We used $(w,r)\in\left\{0.2,0.4,0.6,0.8,1,1.2,1.4,1.6,1.8,2\right\}\otimes\left\{0.55,0.6,0.65,0.7,0.75,0.8,0.85,0.9,0.95,1\right\}$. We ran 20 simulations per parameter set and for 5000 generations each. For each $(w,r)$ pair, we obtained the approximate equilibrium fraction of each strategist as the average over the last 256 generations for each simulation and over the 20 simulations. We used $p\in\{1/5,2/5,3/5,4/5\}$.
@@ III-C Robustness
Figure 8: Robustness of evolutionary outcomes. The baseline case, shown in Fig. 4, uses $p=3/5$, $y_{it}(0)=y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=1/4$, $\beta=10$, $\mu=0$, and $Z=1024$. In the second row of the figure panels, we reduced the initial frequency of $it$ to $y_{it}(0)=1/32$ from $1/4$, while maintaining $y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=\left(1-y_{it}(0)\right)/3$. In the third row, we decreased the selection strength to $\beta=1$ from $10$. In the fourth row, we increased the mutation rate to $\mu=1/Z=1/1024$ from 00. In the fifth row, we decreased the population size to $Z=256$ from $1024$.
The equilibrium of the evolutionary dynamics is robust against variations in parameter values. We verified that order-of-magnitude changes in initial conditions, selection strength, mutation rate or population size produce outcomes qualitatively similar to those of the baseline (see Fig. 8). Notably, the key distinctions in outcomes between the square lattice and heterogeneous networks are well preserved.
@@ IV Discussion
We propose the SNTG and analyse its evolutionary game dynamics. Each player is assumed to play both investor and trustee roles in each generation. Therefore, the mean payoff from both roles drives evolutionary dynamics, unlike the asymmetric NTG in which players have fixed roles (i.e., investor or trustee) and therefore different strategy sets over the evolutionary dynamics [39]. The SNTG reflects scenarios of role-switching. This approach aligns with behavioural experiments of TGs in which participants alternate roles [21]. The original NTG is technically a symmetric game because all players share the same set of strategies [35]. In that game, each player can switch the role as a result of payoff-driven imitation (e.g., an investor turns into an untrustworthy trustee). The original NTG, in which each player has a fixed role at any given time and single-role payoffs drive evolutionary game dynamics, fundamentally differs from the present SNTG, in which each player plays both roles in any generation and the sum of the payoffs from the two roles drives evolutionary dynamics.

### v2

@@ Abstract
Evolutionary game theory, evolutionary dynamics, replicator dynamics, trust game, multiplayer game, symmetry, networks
@@ I-A Evolution of Prosocial Behaviours
Researchers across disciplines have explored how prosocial behaviours evolve among self-interested individuals. A key focus is the evolution of cooperation in social dilemmas, such as the Prisoner’s Dilemma (PD) game and its $N$-player variant, the Public Goods Game (PGG) [1, 2, 3, 4, 5, 6, 7, 8]. In particular, evolutionary game theory enables us to examine how successful strategies proliferate through evolution, i.e., fitness-dependent reproduction and imitation [9, 10]. Evolutionary game theory has also been employed to investigate information dynamics in evolving networks, as well as cooperative packet forwarding in mobile ad hoc networks [11, 12].
Many real-world scenarios involve sequential interactions between two players, as seen in buyer-seller exchanges. Unlike the PD and PGG, which model simultaneous interactions, sequential interactions introduce a trust issue where one player’s decision may leave them vulnerable to exploitation by the other [13]. The concepts of trust and trustworthiness have also gained attention in engineering research [14, 15, 16, 17, 18]. Many problems in these fields are framed as buyer-seller interactions [19]. The Trust Game (TG) is a current standard for formalising non-simultaneous interactions under social dilemmas and is widely used for studying trust and trustworthiness [13, 20, 21, 22, 23, 24]. The TG involves a one-shot sequential interaction between an investor (truster) and a trustee. The binary TG is a simple variant where each role has two strategies: investors can choose to invest or not, and trustees can choose to be trustworthy or untrustworthy (Fig. 1) [22, 25, 26]. Evolutionary game theory predicts that anti-social strategies evolve in the 2-player TG, resulting in reduced payoffs for both players compared to prosocial strategies. Consequently, an additional mechanism is needed for prosocial strategies to evolve in the TG [27, 28, 29]. In practice, interactions often involve multiple participants, so recent attention has been paid to $N$-player TG (NTG) [30, 31, 32, 33, 34, 35], which attempts to generalise the 2-player TG to multiplayer settings, with $N\geq 3$.
Figure 1: Game tree of the asymmetric 2-player binary TG, in which the role of each player is fixed. The payoffs of an investor are shown in green. Those of a trustee are shown in orange. We require $10<r<$ 1, where $r$ represents the relative productivity of the prosocial strategies. Adapted from Ref. [22].
I-B Existing NTG Models and Rationale for Symmetry
@@ I-B2 Asymmetric NTG
Existing NTG models, including the asymmetric NTG detailed previously, typically assume fixed roles with evolutionary dynamics driven by single-role payoffs. However, real-world interactions often involve fluid roles. Even in the standard 2-player TG context, which is inherently asymmetric, symmetrisation (where players alternate between investor and trustee roles) is well-established [42, 29]. This approach reflects the common real-world occurrence of role alternation; for instance, individuals often act as both buyer (investor) and seller (trustee), suggesting that incorporating this dynamic is a worthwhile consideration for modelling trust dynamics.
The relevance of role alternation arguably extends to complex, multi-player settings requiring collective action between groups. Consider, for instance, joint ventures or supply chain clusters where enterprises form a buying consortium (acting as multiple buyers/investors) to jointly buy materials or capacity, needing collective action for scale or access to specialised suppliers. This group of buyers decides whether to collectively engage with a selling alliance (acting as multiple sellers/trustees). Engaging involves placing orders (an act of investment) and implicitly trusting the seller group to act trustworthily, for example, by meeting quality and delivery standards, rather than acting untrustworthily. Crucially, participants often switch roles: the same enterprise in the buying consortium might simultaneously participate in a separate selling alliance (acting as one of multiple sellers/trustees) supplying components to another buying consortium. Such $N$-player scenarios, involving group interactions where entities operate as part of both investor groups and trustee groups, further highlight the potential limitations of fixed-role models. The frequent use of role alternation in behavioural TG experiments also points towards its perceived importance [43, 44, 45].
Given these motivations for considering models with role alternation, a key question arises: how should evolutionary fitness be determined when individuals occupy multiple roles over time? If players frequently switch between being investors and trustees, their overall success likely depends on their performance across both activities. Indeed, existing models of symmetric 2-player TGs often address this by using the average payoff across roles to determine fitness [42, 29]. Evaluating fitness based solely on performance in a single, temporarily held role may therefore be insufficient in dynamic settings. Extending this logic to the multi-player context, we propose a symmetric NTG (SNTG) where evolutionary fitness is explicitly determined by the average payoff accumulated across both the investor and trustee roles. This mechanism contrasts fundamentally with the single-role payoff calculations underpinning fixed-role models or the imitation dynamics in the original NTG.
This paper formally defines this SNTG framework, motivated by the prevalence of role-switching and based on an average payoff structure analogous to that used in symmetric 2-player models, and investigates its evolutionary dynamics. We particularly explore how trust evolves within various population structures and interaction conditions under this symmetric approach, where prosocial outcomes depend on agents successfully navigating both the investor role (e.g., choosing to trust/invest) and the trustee role (e.g., choosing to be trustworthy).
@@ I-C Contributions
Our contributions include:
@@ II-A Strategies
Our SNTG expands upon the asymmetric NTG [34] by symmetrising it. Players alternate between investor and trustee roles, similar to symmetrising 2-player asymmetric games [29, 48, 42]. Investors choose whether to invest or not, while trustees decide to be trustworthy or untrustworthy. A player’s strategy therefore comprises two elements, one for each role: $i$ (i.e., invest) and $n$ (i.e., not to invest) as investor and $t$ (i.e., trustworthy) and $u$ (i.e., untrustworthy) as trustee. There are four possible strategies: $it,iu,nt$, and $nu$.
@@ II-D2 Simulations
Unless we state otherwise, simulations commence with equal proportions of the four strategies ($it$, $iu$, $nt$, and $nu$) uniformly randomly distributed over the $Z=1024$ nodes, which we refer to as the unbiased initial condition. Each simulation runs for 5000 generations. In each generation, the two steps are repeated $Z$ times, allowing each node to update its strategy once per generation on average. For each parameter pair $(w,r)$, we ran 50 independent simulations.
@@ III-A Well-mixed Populations
In a well-mixed population, investment (i.e. trust) does not evolve (Figs. 3 and A.9). Consequently, the average payoff for a player is equal to 00. All trajectories converge to the line of stable equilibria on the NT-NU edge, including the vertex NU (i.e., unanimity of $nu$ players), i.e., $y_{nt}+y_{nu}=1$ with $\frac{r}{r+1}<y_{nu}\leq 1$, of the 3-simplex $\triangle^{3}=\{(y_{it},y_{iu},y_{nt},y_{nu}):y_{it}+y_{iu}+y_{nt}+y_{nu}=1\}$ as the state space (Fig. 3).
Vertices IT, IU and NT of the 3-simplex are unstable equilibria. The entire interior of the IT-NU and IU-NT edges and part of interior of the NT-NU edge given by $0<y_{nu}<\frac{r}{r+1}$ are also unstable equilibria. There is no other equilibria including the interior of the triangles and the tetrahedron. For proofs, see Appendix A-F, A-G, A-H, and A-I. The nonlinearity in the payoff, $w$, has no impact on the evolutionary dynamics. Moreover, the evolutionary outcomes of the SNTG resemble those of the symmetric 2-player TG [29], with both games ultimately resulting in $y_{nt}+y_{nu}=1$, indicating no evolution of investment.
@@ III-B2 Heterogeneous Network
Figure 4: Properties of approximate equilibria of the SNTG in finite networks. (a) Square lattice. (b) Heterogeneous networks. In both (a) and (b), the first four columns show the fraction of each strategy as the function of $w$ and $r$. The fifth column shows the average payoff over all nodes, normalised by that of a population comprised entirely of $it$. We used $(w,r)\in\left\{0.2,0.4,0.6,0.8,1,1.2,1.4,1.6,1.8,2\right\}\otimes\left\{0.55,0.6,0.65,0.7,0.75,0.8,0.85,0.9,0.95,1\right\}$. We ran 50 simulations per parameter set and for 5000 generations each. For each $(w,r)$ pair, we obtained the approximate equilibrium fraction of each strategist as the average over the last 256 generations for each simulation and over the 50 simulations. We used $p\in\{1/5,2/5,3/5,4/5\}$.
@@ III-C Robustness
Figure 8: Robustness of evolutionary outcomes. The baseline case, shown in Fig. 4, uses $p=3/5$, $y_{it}(0)=y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=1/4$, $\beta=10$, $\mu=0$, and $Z=1024$. In the second row of the figure panels, we reduced the initial frequency of $it$ to $y_{it}(0)=1/64$ from $1/4$, while maintaining $y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=\left(1-y_{it}(0)\right)/3$. In the third row, we decreased the selection strength to $\beta=0.1$ from $10$. In the fourth row, we increased the mutation rate to $\mu=0.01$ from 00. In the fifth row, we decreased the population size to $Z=256$ from $1024$. In the sixth row, we increased the mean node degree to $\langle d\rangle=12$ from $4$. For additional results demonstrating robustness, see Figures A.10, A.11, A.12, A.13, and A.14.
The equilibrium of the evolutionary dynamics is robust against variations in parameter values. We demonstrated that order-of-magnitude changes in initial conditions, selection strength, mutation rate, population size or mean node degree produce outcomes qualitatively similar to those of the baseline (see Fig. 8). Notably, the key distinctions in outcomes between the square lattice and heterogeneous networks are well preserved.
@@ IV Discussion
We propose the SNTG and analyse its evolutionary game dynamics. Each player is assumed to play both investor and trustee roles in each generation. Therefore, the mean payoff from both roles drives evolutionary dynamics, unlike the asymmetric NTG in which players have fixed roles (i.e., investor or trustee) and therefore different strategy sets over the evolutionary dynamics [34]. The SNTG reflects scenarios of role-switching. This approach aligns with behavioural experiments of TGs, as well as real-world interactions (such as buyer–seller exchanges), in which participants alternate roles [20]. The original NTG is technically a symmetric game because all players share the same set of strategies [30]. In that game, each player can switch the role as a result of payoff-driven imitation (e.g., an investor turns into an untrustworthy trustee). The original NTG, in which each player has a fixed role at any given time and single-role payoffs drive evolutionary game dynamics, fundamentally differs from the present SNTG, in which each player plays both roles in any generation and the sum of the payoffs from the two roles drives evolutionary dynamics.
@@ A-J Derivation of the threshold $r^{*}$ for the square lattice
A-L Robustness
Figure A.9: Time courses of the SNTG in the infinite well-mixed population with initial conditions in the interior of state space $\triangle^{3}$. The fractions of the different strategies, $y_{it}(t)$, $y_{iu}(t)$, $y_{nt}(t)$, and $y_{nu}(t)$ are shown as a function of time for various initial conditions. (Top row) Unbiased initial state $y_{it}(0)=y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=1/4$. (Bottom) 256 randomly selected initial conditions. We only show $y_{it}(t)$ and $y_{iu}(t)$ for clarity. We observe $y_{it}(t)+y_{iu}(t)\rightarrow 0$ as $t\rightarrow\infty$, regardless of the initial condition; in other words, $y_{nt}(t)+y_{nu}(t)\rightarrow 1$ as $t\rightarrow\infty$. Hence, investment does not evolve. Parameters are the same as those used in Fig. 3.
Figure A.10: Robustness of evolutionary outcomes to intial condition $\left(y_{it}(0),y_{iu}(0),y_{nt}(0),y_{nu}(0)\right)$, where $y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=\frac{1}{3}\left(1-y_{it}(0)\right)$. Changes in the initial condition produce outcomes qualitatively similar to those observed under the baseline condition, $y_{it}(0)=y_{iu}(0)=y_{nt}(0)=y_{nu}(0)=1/4$. This robustness holds for the tested initial conditions which are biased against the most prosocial type, IT (i.e., beginning with a lower initial proportion of IT compared to the baseline). Note that outcomes remain qualitatively similar even when the initial proportion $y_{it}(0)$ of IT is reduced by over an order of magnitude (e.g., from 1/4 to 1/64).
Figure A.11: Robustness of evolutionary outcomes to variations in selection strength, $\beta$. Changes in selection strength yield outcomes that are qualitatively similar to those observed under the baseline condition ($\beta=10$). Note that outcomes remain qualitatively similar even when the selection strength is reduced by two orders of magnitude (e.g., from 10 to 0.1). The range of parameters favourable to the evolution of IT narrows as $\beta$ decreases to 0.1, and then broadens with further decreases in $\beta$, particularly on the square lattice.
Figure A.12: Robustness of evolutionary outcomes to the mutation rate, $\mu$. Variations in the mutation rate yield outcomes qualitatively similar to those observed under the baseline condition of no mutation ($\mu=0$). Note that outcomes remain qualitatively similar even when the mutation rate is increased tangibly from zero (e.g., up to $\mu=0.01$).
Figure A.13: Robustness of evolutionary outcomes to the population size, $Z$. Variations in population size yield outcomes qualitatively similar to those observed under the baseline condition ($Z=1024$). Note that outcomes remain qualitatively similar even when $Z$ is substantially reduced (e.g., from $Z=1024$ to $Z=64$).
Figure A.14: Robustness of evolutionary outcomes to variations in the mean node degree, $\langle d\rangle$. Changes in the mean node degree yield outcomes qualitatively similar to those observed under the baseline condition ($\langle d\rangle=4$). Importantly, the key distinctions in outcomes between the square lattice and heterogeneous networks remain well-preserved across these variations. However, as $\langle d\rangle$ increases, the parameter range conducive to the evolution of IT narrows more markedly in heterogeneous networks compared to square lattices.